a car can go around a curve of radius 43 m at a speed of 12 m/s without skidding. what is the angle of banking?
1
Expert's answer
2021-10-29T11:32:02-0400

"\frac{Nsin\theta}{Ncos\theta}=tan\theta=\frac{mv^2/r}{mg}=\frac{v^2}{gr}"


"\theta=arctan\frac{v^2}{gr}=arctan\frac{12m/s)^2}{(9.81m/s^2)(43m)} =18.85"